OpenStudy (anonymous):

(4e^x)/(e^2x+4e^x+4) integral

OpenStudy (anonymous):

denominator is (e^x+2)^2 go with a u - sub

OpenStudy (anonymous):

How do you factor the denominator? @satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(e^x+2)^2 = e^2x+4e^x+4 u = e^x+2 du = e^x dx integrate(4/u^2 du)